# Deterministic Dynamic Edge-Colouring

### v1

@@ 4.3 Analysing the Dynamic Algorithm
In this subsection, we will analyse the dynamic algorithm above. We begin by analysing how precise the estimates $\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v)$ in the relevant data structures are. Then we show that the algorithm actually maintains Invariant 20, thus establishing correctness via Theorem 21. Then we analyse the recourse of the algorithm, before we pick explicit implementations of the data structures and analyse the final running time.
@@ 4.3.1 Analysis of Round-Robin
In this section, we will analyse the round-robin scheme applied by the algorithm. In particular, we will show that the algorithm maintains Invariant 20. We will do this by showing that even-though the algorithm sometimes works with (slightly) out-dated information, the information is still precise enough for the algorithm to detect necessary changes and perform them efficiently.
To this end, we first show that the data structures used by the algorithm contain good estimates. Before doing this, we will introduce some notation. For any data structure $\mathcal{D}$, we will let $(\mathcal{D})_{\tau}$ be the state of the data structure after the $(\tau-1)^{th}$ call to ReColour concluded, and just before the $\tau^{th}$ call to ReColour commences. Similarly, for any value $p(x)$ we will let $(\tilde{p}(x))_{\mathcal{D},\tau}$ denote the estimate of this value stored in data-structure $\mathcal{D}$ at time $\tau$, and similarly we will let $(p(x))_{\tau}$ denote the actual value of the parameter at time $\tau$. If the data-structure is clear from the context, we will leave out this argument to ease readability.
Lemma 24.
@@
    $(\tilde{E}(u,uv,\kappa))_{\tau}=((\tilde{s}_{c(uv)}(u))+(\tilde{s}_{c(uv)}(v)))-((\tilde{s}_{\kappa}(u))+(\tilde{s}_{\kappa}(v)))_{\tau}$
@@
    $(E(u,uv,\kappa))_{\tau}=((s_{c(uv)}(u))+(s_{c(uv)}(v)))-((s_{\kappa}(u))+(s_{\kappa}(v)))_{\tau}$
Proof.
We will show item 1)1)1 ) first. Let $\tau^{\prime}<\tau$ be the last time before time $\tau$, where $\tilde{S}$ was updated. We then have that
$(\tilde{S})_{\tau}=(\tilde{S})_{\tau^{\prime}}=(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau^{\prime}}=((s_{\kappa}(u))_{\tau^{\prime}}+(s_{\kappa}(v))_{\tau^{\prime}}$
and so it is sufficient to estimate differences of the form $|s_{\kappa}(w)_{\tau}-s_{\kappa}(w))_{\tau^{\prime}}|$. To this end, we note that due to the round-robin scheme being run every time an edge incident to $w$ changes its colour, we can change $s_{\kappa}(w)$ by at most
@@
$|(\tilde{S})_{\tau}-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})|$ $\leq|((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})|$
$\leq|(s_{\kappa}(u))_{\tau}-(s_{\kappa}(u))_{\tau}|+|(s_{\kappa}(v))_{\tau}-(s_{\kappa}(v))_{\tau}|$
@@
$(\tilde{E}(u,uv,\kappa))_{\tau}=(\tilde{s}_{\kappa^{\prime}}(u)+\tilde{s}_{\kappa^{\prime}}(v))_{\tau}-(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau}$
@@
Thus we can apply $1$ to conclude that:
$|\displaystyle|(E(u,uv,\kappa)_{\tau}-(\tilde{E}(u,uv,\kappa))_{\tau}|$$\leq|(\tilde{s}_{\kappa^{\prime}}(u)+\tilde{s}_{\kappa^{\prime}}(v))_{\tau}-((s_{\kappa^{\prime}}(u))_{\tau}+(s_{\kappa^{\prime}}(v))_{\tau})|$
$+|((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})-(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau}|$
@@
Next we will argue that this information is accurate enough to maintain Invariant 20, provided that the algorithm terminates. Afterwards, we show that the algorithm always terminates and in fact does so with small amortised recourse.
Lemma 25.
Let $G$ be a dynamic graph, and let $c_{i}$ be the colouring in $G_{i}$ maintained by the algorithm after $i$ updates. Assume that the algorithm terminates after each update and that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then $c_{i}$ satisfies Invariant 20 for all $i$.
Proof.
It suffices to show that after every update, we must have $\Delta s_{\kappa}(w)<\eta$ for all vertices $w$ and all colours $\kappa$. We first note that by Lemma 24, it follows that if $\Delta\tilde{s}_{\kappa}(w)\leq\tfrac{\eta}{2}$, then we have $\Delta s_{\kappa}(w)\leq\Delta\tilde{s}_{\kappa}(w)+\tfrac{\eta}{8}<\eta$.
@@
[…]
To this end, we note that after a call to Initialize-DS the degree of $w$ is exactly one, and so since $\eta\geq 10$, this call can never cause $\Delta\tilde{s}_{\kappa}(w)>\tfrac{\eta}{2}$.
[…]
@@ 4.3.2 Analysing the Recourse
In this sub-section, we analyse the recourse of the algorithm and show that the it terminates. We first show that any edge re-coloured by the algorithm results in a significant lowering of the surplus of this edge. This allows us to show that a certain global potential of the graph drops enough to ensure that the amortised number of recolourings become bounded.
To this end, we first argue that any recolouring performed by the algorithm lowers the surplus of an edge significantly:
Lemma 26.
@@
Proof.
Any $\kappa$-coloured edge incident to $u$ or $v$ is only recoloured during time-step $\tau$ if $\Delta\tilde{s}_{\kappa}(u)\geq\tfrac{\eta}{2}$ or $\Delta\tilde{s}_{\kappa}(v)\geq\tfrac{\eta}{2}$. Since $e$ was chosen as an edge maximising the surplus difference, it must be the case that
$\tilde{\delta}^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}$ $:=((\tilde{s}_{\kappa_{1}}(u)+\tilde{s}_{\kappa_{1}}(v))-(\tilde{s}_{\kappa_{2}}(u)+\tilde{s}_{\kappa_{2}}(v)))_{\tau-1}$
@@
where the second equality follows from the fact that $MSD_{\kappa}(u)$ always is updated by a difference of surpluses stored in some BBST $T^{s}(uv)$. Hence, it follows that:
@@
Consequently, we can apply Lemma 24 part 1) in order to observe that:
@@
As we will show below, Lemma 26 implies that every recolouring performed by the algorithm will lower the potential substantially. In particular, this ensures that the amortised recourse of the dynamic algorithm is low. This is shown in the following Lemma:
Lemma 27.
Let $\tilde{d}=\max_{c}\max_{v}\tilde{s}(v)$ be the maximum surplus of any colouring satisfying Invariant 20 with $10<\varepsilon<$ 1. Then the amortised recourse of the algorithm is upper bounded by $16\frac{\tilde{d}}{\eta}\leq\frac{2^{10}\log n}{\varepsilon^{2}}=O(\tfrac{\log n}{\varepsilon^{2}})$ provided that $\eta=\lfloor\tfrac{\varepsilon^{2}}{128\log n}\rfloor\tfrac{\Delta_{\max}}{t}\geq 16$ and $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$.
@@
Proof.
We first note that after inserting and colouring any edge $e=uv$, the potential can increase by at most $2(\tilde{d}+1)$, since between insertions the algorithm guarantees that Invariant 20 holds, provided that the algorithm terminates, by Lemma 25. Note that a deletion will always lower the potential.
The second observation is that each recolouring will lower the potential by $\Omega(\eta)$. Indeed, if $e$ is recoloured from having colour $\kappa_{1}$ to having colour $\kappa_{2}$ during time-step $\tau$, then it follows from Lemma 26 that:
@@
Indeed, the above holds since recolouring $e$ reduces $s_{\kappa_{1}}(u)$ and $s_{\kappa_{1}}(v)$ by 1, and hence $\Phi$ by $(s_{\kappa_{1}}(u))_{\tau-1}+(s_{\kappa_{1}}(v))_{\tau-1}$. On the other hand, $s_{\kappa_{2}}(u)$ and $s_{\kappa_{2}}(v)$ are both increased by one, which increases the potential by $(s_{\kappa_{2}}(u)))_{\tau-1}+(s_{\kappa_{2}}(v)))_{\tau-1}+2$.
Since $\Phi$ is always non-negative, since deletions do not lower $\Phi$, and since insertions increase $\Phi$ by at most $2(\tilde{d}+1)$, we find that we can witness at most:
@@
recolourings over a sequence of updates containing $r$ insertions. Consequently, the algorithm always terminates and it does so with recourse bounded by $O(\frac{\tilde{d}}{\eta})$. To recover the last part of the Lemma, note that Theorem 21 gives that $\tilde{d}\leq(1+\varepsilon)\tfrac{\Delta_{\max}}{t}$ and that $\eta=\lfloor\tfrac{\varepsilon^{2}}{128\log n}\rfloor\tfrac{\Delta_{\max}}{t}$. ∎
We briefly remark that it is quite important that $\eta\geq 16$. This ensures that any recolouring actually lowers the potential, and that one does not end up in a situation where the algorithm continually recolours the same edges. In the next section, we will analyse the update time as a function of the number of recolourings i.e. the recourse performed by the algorithm. We will then use Lemma 27 to upper bound the total update time.
@@ 4.3.3 Analysing the Update Time
[…]
  • Not counting the recursive call, for all edges $e$, we can perform one call to $\emph{{ReColour}}(e)$ in $O(t^{3}\varepsilon^{-2}\log^{2}m)$ time.
@@
[…]
Proof.
We will only require our BBST’s to support very basic operations, so any standard BBST such as red-black trees [CLRS22] will do.
@@
For the second item, we note that each update to $T^{s}_{uv}$ takes $O(\log t)$ time, since we can compute $s_{\kappa}(u)+s_{\kappa}(v)$ in constant time. Each update to $MSD_{\kappa}(v)$ can also be performed in $O(\log n)$ time. Indeed, we can compute $(s_{c(uv)}(u)+s_{c(uv)}(v))-(s_{\kappa}(u)+s_{\kappa}(v))$ in $O(1)$ time, and insert it in $O(\log m)$ time. Similarly, by storing a pointer to the element at $w$, we can also remove the old entrance in $O(\log m)$ time. All of the above has to be performed $t$ times, leaving us with the stated update time.
[…]
@@
For the sixth and final item, we note that extracting the maximum entrance in a BBST can be performed in $O(\log n)$ time by a standard search always going to the largest child. ∎
Everytime we call ReColour on some edge, we perform a recolouring. Hence, by Lemma 27, after $i$ insertions, we have performed at most $16\cdot{}i\cdot{}\frac{\tilde{d}}{\eta}$ calls to ReColour. Thus, in an amortised sense, the number of calls to ReColour is $O(\frac{\tilde{d}}{\eta})$ per update. If we ignore the recursive calls, then by Lemma 28 the time spent for each call is dominated by the call to Round-Robin, and so each call to Recolour takes only $O(t^{3}\varepsilon^{-2}\log^{2}m)$ time. Hence, we have shown the following bound on the amortised running time of insert and delete.
Lemma 29.
@@
Combining Lemma 29, Lemma 27, and Lemma 25 now gives Theorem 3.

### v2

@@ 4.3 Analysing the Dynamic Algorithm
In this subsection, we will analyse the dynamic algorithm above. We begin by analysing how precise the estimates $\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v)$ in the relevant data structures are. Then we show that the algorithm actually maintains Invariant 4.2, thus establishing correctness via Theorem 4.3. Then we analyse the recourse of the algorithm, before we pick explicit implementations of the data structures and analyse the final running time.
@@ 4.3.1 Analysis of Round-Robin
In this section, we will analyse the round-robin scheme applied by the algorithm. In particular, we will show that the algorithm maintains Invariant 4.2. We will do this by showing that even-though the algorithm sometimes works with (slightly) out-dated information, the information is still precise enough for the algorithm to detect necessary changes and perform them efficiently.
To this end, we first show that the data structures used by the algorithm contain good estimates. Before doing this, we will introduce some notation. For any data structure $\mathcal{D}$, we will let $(\mathcal{D})_{\tau}$ be the state of the data structure after the $\tau^{th}$ call to RoundRobin concluded, and just before the $(\tau+1)^{th}$ call to RoundRobin commences. Similarly, for any value $p(x)$ we will let $(\tilde{p}(x))_{\mathcal{D},\tau}$ denote the estimate of this value stored in data structure $\mathcal{D}$ at time $\tau$, and we will let $(p(x))_{\tau}$ denote the actual value of the parameter at time $\tau$. We will say that something happens during time step $\tau$ if it happens between the $(\tau-1)^{th}$ and $\tau^{th}$ call to RoundRobin. If the data structure is clear from the context, we will leave out this argument to ease readability.
Lemma 4.6.
@@
    $(\tilde{E}(u,uv,\kappa))_{\tau}=((\tilde{s}_{c(uv)}(u)+\tilde{s}_{c(uv)}(v))-(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v)))_{\tau}$
@@
    $(E(u,uv,\kappa))_{\tau}=((s_{c(uv)}(u)+s_{c(uv)}(v))-(s_{\kappa}(u)+s_{\kappa}(v)))_{\tau}$
Proof 8.
We will show item $1)$ first. Let $\tau^{\prime}\leq\tau$ be the last time before (or at) time $\tau$, where $\tilde{S}$ was updated. We then have that
$(\tilde{S})_{\tau}=(\tilde{S})_{\tau^{\prime}}=(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau^{\prime}}=(s_{\kappa}(u))_{\tau^{\prime}}+(s_{\kappa}(v))_{\tau^{\prime}}$
and so it is sufficient to estimate differences of the form $|(s_{\kappa}(w))_{\tau}-(s_{\kappa}(w))_{\tau^{\prime}}|$. To this end, we note that due to the round-robin scheme being run every time an edge incident to $w$ changes its colour, we can change $s_{\kappa}(w)$ by at most
@@
$|(\tilde{S})_{\tau}-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})|$ $\leq|((s_{\kappa}(u))_{\tau^{\prime}}+(s_{\kappa}(v))_{\tau^{\prime}})-((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})|$
$\leq|(s_{\kappa}(u))_{\tau^{\prime}}-(s_{\kappa}(u))_{\tau}|+|(s_{\kappa}(v))_{\tau^{\prime}}-(s_{\kappa}(v))_{\tau}|$
@@
$(\tilde{E}(u,uv,\kappa))_{\tau}=(\tilde{s}_{\kappa^{\prime}}(u)+\tilde{s}_{\kappa^{\prime}}(v))_{\tau^{\prime}}-(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(v))_{\tau^{\prime}}$
for some $\tau^{\prime}\leq\tau$ denoting the last time before (or at) $\tau$ that $MSD_{\kappa}(u)$ was updated. Observe that any time we update the entry $\tilde{E}(u,uv,\kappa)$ in $MSD_{\kappa}(u)$, we do so right after updating both $\tilde{s}(u)_{\kappa^{\prime}}$ and $\tilde{s}(v)_{\kappa^{\prime}}$. Therefore, we can in fact write:
$(\tilde{E}(u,uv,\kappa))_{\tau}=(s_{\kappa^{\prime}}(u)+s_{\kappa^{\prime}}(v))_{\tau^{\prime}}-(s_{\kappa}(u)+s_{\kappa}(v))_{\tau^{\prime}}$
@@
Thus similarly to in case $1.$, we conclude that:
$|(E(u,uv,\kappa)_{\tau}-(\tilde{E}(u,uv,\kappa))_{\tau}|$ $\leq|(s_{\kappa^{\prime}}(u)+s_{\kappa^{\prime}}(v))_{\tau^{\prime}}-((s_{\kappa^{\prime}}(u))_{\tau}+(s_{\kappa^{\prime}}(v))_{\tau})|$
$+|((s_{\kappa}(u))_{\tau}+(s_{\kappa}(v))_{\tau})-(s_{\kappa}(u)+s_{\kappa}(v))_{\tau^{\prime}}|$
@@
Next we will argue that this information is accurate enough to maintain Invariant 4.2, provided that the algorithm terminates. Afterwards, we show that the algorithm always terminates and in fact does so with small amortised recourse.
Lemma 4.7.
Let $G$ be a dynamic graph, and let $c_{i}$ be the colouring in $G_{i}$ maintained by the algorithm after $i$ updates. Assume that the algorithm terminates after each update and that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then $c_{i}$ satisfies Invariant 4.2 for all $i$.
Proof 9.
It suffices to show that after every update, we must have $\Delta s_{\kappa}(w)<\eta$ for all vertices $w$ and all colours $\kappa$. We first note that by Lemma 4.6, it follows that if $\Delta\tilde{s}_{\kappa}(w)\leq\tfrac{\eta}{2}$, then we have $\Delta s_{\kappa}(w)\leq\Delta\tilde{s}_{\kappa}(w)+\tfrac{\eta}{8}<\eta$.
@@
Since the round-robin scheme, which updates the values stored in $MSD_{\kappa}(w)$, is only called on vertices $w$ either directly incident to edges that were inserted, deleted or recoloured or vertices neighbouring such vertices, the desired statement holds for any other vertex by induction. In particular, $MSD_{\kappa}(w)$ is only changed after deleting an edge, giving an edge a new colour, or during a call to Update-DS. The latter only happens during a call to $\texttt{Round-Robin}(uv)$ or imedditaly after recolouring an edge.
We first note that we need only worry about vertices directly incident to inserted, deleted or recoloured edges, and not neighbours of such vertices. Indeed, we always update an edge $uv$’s contribution to $MSD_{\kappa}(w)$ for $w\in\{u,v\}$ for both endpoints to the same value. This means that after a call to $\texttt{Update-DS}(uv)$ or a direct update terminates, we always have $\tilde{E}(u,uv,\kappa)=\tilde{E}(v,uv,\kappa)$ i.e. for all $\tau$ we have $(\tilde{E}(u,uv,\kappa))_{\tau}=(\tilde{E}(v,uv,\kappa))_{\tau}$. Hence, it is sufficient to verify that any endpoint $w$ directly incident to an insertion, deletion or recolouring satisfies $\Delta\tilde{s}_{\kappa}(w)\leq\tfrac{\eta}{2}$ after the update procedure ends.
After every $\texttt{Round-Robin}(uv)$ call or direct update, any vertex $w$ with $\Delta\tilde{s}_{\kappa}(w)>\tfrac{\eta}{2}$ is immediately added to a queue $Q$. The vertex $w$ can only leave $Q$ in two ways: either because it is verified at some time step $\tau$ that $(\Delta\tilde{s}_{\kappa}(w))_{\tau}\leq\tfrac{\eta}{2}$ or because some $\kappa$-coloured edge incident to $w$ is recoloured. In the first case, $(w,\kappa)$ can only enter $Q$ again if some edge incident to it is recoloured at some time-step $\tau^{\prime}>\tau$ in which case the exact same argument can be applied again. If no such future update occurs, we immediately have that $\Delta\tilde{s}_{\kappa}(w)\leq\tfrac{\eta}{2}$ after the update procedure ends. In the latter case, it is checked after the recolouring whether $\Delta\tilde{s}_{\kappa}(w)>\tfrac{\eta}{2}$ and if so, then $w$ is added back to $Q$ but at a later time-step. If $\Delta\tilde{s}_{\kappa}(w)\leq\tfrac{\eta}{2}$, then $w$ immediately leaves $Q$ and similar arguments to the first case applies.
Since we have assumed that the algorithm terminates, we must eventually end up in the first case.
@@ 4.3.2 Analysing the Recourse
In this subsection, we analyse the recourse of the algorithm and show that the it terminates. We first show that any edge re-coloured by the algorithm results in a significant lowering of the surplus of this edge. This allows us to show that a certain global potential of the graph drops enough to ensure that the amortised number of recolourings becomes bounded.
Lemma 4.8.
@@
Proof 10.
Any $\kappa$-coloured edge incident to $u$ or $v$ is only recoloured during time-step $\tau$ if $(\Delta\tilde{s}_{\kappa}(u))_{\tau-1}\geq\tfrac{\eta}{2}$ or $(\Delta\tilde{s}_{\kappa}(v))_{\tau-1}\geq\tfrac{\eta}{2}$. Since $e$ was chosen as an edge maximising the surplus difference, it must be the case that
$\tilde{\delta}^{uv}_{\kappa_{1}\rightarrow\kappa_{2}}$ $:=(\tilde{s}_{\kappa_{1}}(u)+\tilde{s}_{\kappa_{1}}(v)-(\tilde{s}_{\kappa_{2}}(u)+\tilde{s}_{\kappa_{2}}(v)))_{\tau-1}$
@@
where the second equality follows from the fact that $MSD_{\kappa}(u)$ always is updated by a difference of surpluses stored in BBSTs of of the form $T^{s}(uv)$ immediately after all of these data structures have been updated. Hence, it follows that:
@@
Consequently, we can apply Lemma 4.6 part 1) in order to observe that:
@@
As we will show below, Lemma 4.8 implies that every recolouring performed by the algorithm will lower the potential substantially. In particular, this ensures that the amortised recourse of the dynamic algorithm is low. This is shown in the following Lemma:
Lemma 4.9.
Let $\tilde{d}=\max_{c}\max_{v}\tilde{s}(v)$ be the maximum surplus of any colouring satisfying Invariant 4.2 with $0<\varepsilon<1$. Then the amortised recourse of the algorithm is upper bounded by $16\frac{\tilde{d}}{\eta}\leq\frac{2^{10}\log n}{\varepsilon^{2}}=O(\tfrac{\log n}{\varepsilon^{2}})$ provided that $\eta=\lfloor\tfrac{\varepsilon^{2}}{128\log n}\rfloor\tfrac{\Delta_{\max}}{t}\geq 16$ and $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$.
@@
Proof 11.
We first note that after inserting and colouring any edge $e=uv$, the potential can increase by at most $2(\tilde{d}+1)$, since between insertions the algorithm guarantees that Invariant 4.2 holds, provided that the algorithm terminates, by Lemma 4.7. Note that a deletion will always lower the potential.
The second observation is that each recolouring will lower the potential by $\Omega(\eta)$. Indeed, if $e$ is recoloured from having colour $\kappa_{1}$ to having colour $\kappa_{2}$ during time-step $\tau$, then it follows from Lemma 4.8 that:
@@
Indeed, the above holds since recolouring $e$ reduces $s_{\kappa_{1}}(u)$ and $s_{\kappa_{1}}(v)$ by 1, and hence $\Phi$ by $(s_{\kappa_{1}}(u))_{\tau-1}+(s_{\kappa_{1}}(v))_{\tau-1}$. On the other hand, $s_{\kappa_{2}}(u)$ and $s_{\kappa_{2}}(v)$ are both increased by one, which increases the potential by $(s_{\kappa_{2}}(u)))_{\tau-1}+(s_{\kappa_{2}}(v)))_{\tau-1}+2$. Note also that at most one edge can be recoloured during a time-step.
Since $\Phi$ is always non-negative, deletions do not lower $\Phi$, and since insertions increase $\Phi$ by at most $2(\tilde{d}+1)$, we find that we can witness at most:
@@
recolourings over a sequence of updates containing $r$ insertions. Consequently, the algorithm always terminates and it does so with recourse bounded by $O(\frac{\tilde{d}}{\eta})$. To recover the last part of the Lemma, note that Theorem 4.3 gives that $\tilde{d}\leq(1+\varepsilon)\tfrac{\Delta_{\max}}{t}$ when $\eta=\lfloor\tfrac{\varepsilon^{2}}{128\log n}\rfloor\tfrac{\Delta_{\max}}{t}$.
We briefly remark that it is quite important that $\eta\geq 16$. This ensures that any recolouring actually lowers the potential, and that one does not end up in a situation where the algorithm continually recolours the same edges. In the next section, we will analyse the update time as a function of the number of recolourings i.e. the recourse performed by the algorithm. We will then use Lemma 4.9 to upper bound the total update time.
@@ 4.3.3 Analysing the Update Time
  • For all edges $uv$ incident to $v$, we can perform $\emph{{Initialize-DS}}(uv)$ in $O(t\log m)$ time.
  • For all edges $uv$ incident to $v$, we can perform $\emph{{Update-DS}}(uv)$ in $O(t\log m)$ time.
  • For all edges $uv$ incident to $v$, we can perform $\emph{{Round-Robin}}(uv)$ in $O(t^{3}\varepsilon^{-2}\log^{2}m)$ time.
  • We can perform one call to $\emph{{ReColour}}(e)$ in $O(t^{3}\varepsilon^{-2}\log^{2}m)$ time.
@@
  • We can perform $\emph{{ExtractMinColour}}(T^{s}_{uv})$ in $O(\log m)$ time.
Proof 12.
We will only require our BBST’s to support very basic operations, so any standard BBST such as red-black trees [cormen2022introduction] will do.
@@
For the second item, we note that each update to $T^{s}_{uv}$ takes $O(\log t)$ time, since we can compute $s_{\kappa}(u)+s_{\kappa}(v)$ in constant time. Each update to $MSD_{\kappa}(v)$ can also be performed in $O(\log n)$ time. Indeed, we can compute $(s_{c(uv)}(u)+s_{c(uv)}(v))-(s_{\kappa}(u)+s_{\kappa}(v))$ in $O(1)$ time, and insert it in $O(\log m)$ time. Similarly, by storing pointers to the elements in a BBST at $w$, we can also remove the old entrance in $O(\log m)$ time. All of the above has to be performed $t$ times, leaving us with the stated update time.
For the third item, we note that we perform $2t\cdot{}\lceil 500\tfrac{\Delta}{\eta}\rceil$ calls to $\texttt{Update-DS}(e)$, which dominates the update time. Indeed, we can index into the doubly-linked-list in $O(1)$ time, provided that we have stored a pointer to the element, we wish to index, as well as its position in the linked list. Applying the second item therefore yields the running time.
For the fourth item, we note that we can perform $\texttt{ExtractMinColour}(T^{s}_{uv})$ in $\log t$ time by simply computing the minimum in the BBST. Hence, if we disregard further recursive call, the update time is dominated by the call to $\texttt{Round-Robin}(uv)$, so the statement follows from the third item.
@@
For the sixth and final item, we note that extracting the maximum entrance in a BBST can be performed in $O(\log n)$ time by a standard search always going to the largest child.
Everytime we call ReColour on some edge, we perform a recolouring. Hence, by Lemma 4.9, after $i$ insertions, we have performed at most $16\cdot{}i\cdot{}\frac{\tilde{d}}{\eta}$ calls to ReColour. Thus, in an amortised sense, the number of calls to ReColour is $O(\frac{\tilde{d}}{\eta})$ per update. If we ignore the recursive calls, then by Lemma 4.10 the time spent for each call is dominated by the call to Round-Robin, and so each call to Recolour takes only $O(t^{3}\varepsilon^{-2}\log^{2}m)$ time. Each recolouring puts at most $2t$ pairs unto $Q$. The time required to process such pairs which do not result in a recolouring can be, in an amortised sense, ascribed to the recolouring that pushed the pair unto $Q$. This time is also dominated by the time is dominated by the call to Round-Robin.
Hence, we have shown the following bound on the amortised running time of insert and delete.
Lemma 4.11.
@@
Combining Lemma 4.7, Lemma 4.9, and Lemma 4.11 now gives Theorem 1.3.